Q8.

Calculate the standard free energy change for the combustion of methane at 25oC. Assume that all reactants and products are gases. Provide an answer correct to 3 significant figures. Do not use scientific notation or include units to report your answer.


1
Expert's answer
2022-02-14T17:04:04-0500

"CH_{4(g)}+2O_{2(g)}\to\>CO_{2(g)}+2H_2O_{(g)}"


"C_{(s)}+2H_{2(g)}\to\>CH_{4(g)}" "\Delta\>H_f^\theta=-74.9kJ/mol"


"C_{(s)}+O_{2(g)}\to\>CO_{2(g)}\>\>\>\Delta\>H_f^\theta=-393.5kJ/mol"



"2H_{2(g)}+O_{2(g)}\to\>2H_2O_{(g)}\>\>\Delta\>H_f^\theta=-241.8kJ/mol"



"\Delta\>H_{comb}^\theta(CH_4)=\Delta\>H_f^\theta(CO_2)+2\Delta\>H_f^\theta(H_2O)-\Delta\>H_f^\theta(CH_4)"


"=-393.5+2(-241.8)-(-74.9)"

"=-802.2kJ\/mol"

"=-802"